OpenStudy (anonymous):

outcome HH-28 # times tossed HT -22 TH-34 TT-16 1. experimental probability is 3% greater than the theoretical probabilty? or 2. The theoretical probability is 3% greater than the experimental probability? or is it 3. that they are both equal? @eSpeX

OpenStudy (anonymous):

I think equal?

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

Can you show me your work?

OpenStudy (anonymous):

I don't know if you remember me? My computer wouldn't let me upload took me 20 minutes. I need new computer : (

OpenStudy (whpalmer4):

Just type it. how do you find the experimental probability, in general?

OpenStudy (anonymous):

one time you got it one your Ipad

OpenStudy (anonymous):

That is where I need help.

OpenStudy (anonymous):

I want to really understand it.

OpenStudy (whpalmer4):

Okay. We're trying to find the probability of tossing HH?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I also have been doing what you said practicing with coins

OpenStudy (whpalmer4):

Okay. So it looks like we did 100 trials, and got the following results: HH 28 times out of 100 HT 22 times out of 100 TH 34 times out of 100 TT 16 times out of 100 Our experimental probability for each outcome is simply the number of times we got that outcome/number of trials. For TT, for example, P_exp would be 16/100 = 0.16 or 16%.

OpenStudy (whpalmer4):

Clear so far?

OpenStudy (anonymous):

looking at it just sec

OpenStudy (anonymous):

ohh ok yes yes!

OpenStudy (whpalmer4):

Now, do you know how to find the theoretical probability for each of those combinations?

OpenStudy (anonymous):

divide?

OpenStudy (whpalmer4):

Well, wouldn't one expect them all to be equal? Here are the four possible outcomes: H H H T T T T H with a truly fair coin, each of those has an equal probability, right?

OpenStudy (anonymous):

Yes : ) and that is what I was thinking but wasn't 100%

OpenStudy (whpalmer4):

There's exactly 1 way to flip Heads followed by Heads. There are 4 possible outcomes, so the theoretical probability of Heads followed by Heads is 1/4 or 0.25 or 25%. Similarly the other 3 are each the same. So, with that in mind, would you like to revisit your earlier answer? :-)

OpenStudy (anonymous):

They are equal : )

OpenStudy (whpalmer4):

No.... ;-(

OpenStudy (anonymous):

: (

OpenStudy (whpalmer4):

The 4 theoretical probabilities are equal to each other, yes, but the question asks you to compare the theoretical and experimental probabilities of flipping Heads followed by Heads, no?

OpenStudy (anonymous):

So the experimental is 3% greater?

OpenStudy (whpalmer4):

Experimental probability for HH was 28%, and theoretical probability was 25%. 28%-25%=3% so experimental probability is 3% greater.

OpenStudy (anonymous):

yes ok I'm slowly getting it.

OpenStudy (anonymous):

How did you get 25% ? I know dumb question.

OpenStudy (whpalmer4):

1/4 = 0.25, to convert a decimal to a percentage, multiply by 100%, 0.25*100% = 25%

OpenStudy (whpalmer4):

Or are you asking how I figured out that it was 1/4?

OpenStudy (anonymous):

ok because there are 4 outcomes shown and HH-28 ok got it

OpenStudy (anonymous):

Figuring it out so if I had another problem like it I can see how you solved it.

OpenStudy (whpalmer4):

Right. The 4 possible outcomes are HH HT TH TT, so there is 1/4 chance theoretically of getting HH: 1 of the 4 possible outcomes is HH. Alternatively, you could think of it by multiplying probabilities: 1/2 for the first coin flip to give you a heads result, and 1/2 for the second coin flip to give you a heads result, so 1/2 of the time * 1/2 of the time you get 2 in a row, and 1/2 * 1/2 = 1/4.

OpenStudy (anonymous):

I have another question. I will close this first and reopen a new question?

OpenStudy (whpalmer4):

Okay.

OpenStudy (anonymous):

Ok I just need to know how to figure it. just sec